OpenStudy (anonymous):

Check my answer please Suppose that a restaurant chain claims that its bottles of ketchup contain 24 ounces of ketchup on average, with a standard deviation of 0.8 ounces. If you took a sample of 49 bottles of ketchup, what would be the approximate 95% confidence interval for the mean number of ounces of ketchup per bottle in the sample? A. 24 ± 0.029 B. 24 ± 0.057 C. 24 ± 0.114 D. 24 ± 0.229 I think it's d

OpenStudy (anonymous):

@SolomonZelman

OpenStudy (anonymous):

@jim_thompson5910

OpenStudy (anonymous):

@perl

jimthompson5910 (jim_thompson5910):

yeah you should have ME = t*s/sqrt(n) ME = 2.0106*0.8/sqrt(49) ME = 0.22978 due to rounding error, the margin of error is a bit off, but it's close enough to 0.229

OpenStudy (anonymous):

Thank you